# Data-Driven Nonconvex Reachability Analysis using Exact Multiplication

### v1

@@ I Introduction
Data-driven approaches for reachability can be categorized into deterministic frameworks with bounded uncertainties and stochastic frameworks with probabilistic uncertainty models. Within the deterministic framework, several set representations have been employed to obtain overapproximated reachable sets. For instance, [1] employs matrix zonotopes for representing the set of possible models consistent with noisy data, which is then used to approximate the reachable set as a constrained zonotope, a set representation introduced by [2]. To incorporate prior knowledge and reduce conservatism, [3] employs constrained matrix zonotopes and extends the method to polynomial systems and Lipschitz nonlinear systems. Similarly, [4] uses linear time-varying approximations with bounded error terms for nonlinear systems and proposes an ellipsoidal over-approximation of the reachable set by solving a convex optimization problem. For hybrid or logical systems, [5] and [6] introduce hybrid zonotopes and logical polynomial zonotopes, respectively, for representing reachable sets.
@@
Within the stochastic framework, [7, 8] employs kernel distribution embeddings within reproducing kernel Hilbert spaces for data-driven reachability of stochastic systems. Alternatively, [9] uses empirical inverse Christoffel functions with probably approximately correct (PAC) guarantees, whereas [10] employs Gaussian process state space models for exact probability measures of state trajectories. A specific sub-category within this framework focuses on data-driven reachability analysis for human-in-the-loop systems. In this regard, [11] and [12] employ Gaussian mixture models and empirically optimized disturbance bounds, respectively.
@@
Deterministic approaches to data-driven reachability are preferred when constraint violations are unacceptable, e.g., collision avoidance. This is why regulatory frameworks for safety-critical systems (e.g., DO-178C: Software Considerations in Airborne Systems and Equipment Certification) often require deterministic guarantees rather than probabilistic ones. Furthermore, when data is limited or uncertainty distributions are unknown, deterministic approaches avoid potentially incorrect distributional assumptions, leading to robust guarantees even against worst-case scenarios.
@@
Handling non-convex sets in deterministic reachability analysis is quite challenging. While convex representations have a computational advantage, they sacrifice accuracy by overapproximating the reachable set [3]. Recent works [13, 14, 15] have addressed non-convex reachable sets to some degree. However, accurately representing and propagating non-convex sets without introducing excessive conservatism or computational complexity remains an open problem.
@@
The challenge of non-convexity in data-driven reachability is particularly evident in linear systems where one obtains a reachable set by multiplying a set of system matrices consistent with data with a set of possible states. For instance, in [3], reachable sets and a set of possible system matrices consistent with data are described using constrained zonotopes (CZs) and constrained matrix zonotopes (CMZs), respectively. However, CZs are not closed under multiplication with CMZs. This is the reason why convex set-based approaches to reachability typically resort to approximations that overbound the resulting set, leading to conservatism.
@@
In this paper, we address this challenge in deterministic data-driven reachability by employing constrained polynomial zonotopes (CPZs) recently introduced by [16]. While [17] presents an exact multiplication for matrix zonotopes (MZs) and polynomial zonotopes (PZs), it does not incorporate constraints into the multiplication itself. Therefore, our key contribution is to add the constraints of the sets in the exact multiplication operation by leveraging the closure property of CPZs when multiplied with CMZs. This extension preserves the non-convex geometry without introducing conservatism, thereby enabling more accurate safety verification for systems that can be more precisely characterized as CMZs—as opposed to just MZs—when dealing with non-convex reachable sets.
@@ II-B Problem Statement
Our goal is to estimate the exact reachable set of the system (1) using input-state data, without explicit knowledge of the system matrices $\Phi_{\text{tr}}$ and $\Gamma_{\text{tr}}$. In particular, we are interested in the case where the initial set $\mathcal{X}_{0}$ is non-convex, and we must compute a reachable set that preserves this non-convex geometry without resorting to convex relaxations.
@@ II-C Set Representations
Given an offset $c\!\in\!\mathbb{R}^{n}$, generator $G\!\in\!\mathbb{R}^{n\times h}$, and constraints $A\in\mathbb{R}^{n_{c}\times h}$ and $b\in\mathbb{R}^{n_{c}}$, a constrained zonotope (CZ) is a compact set in $\mathbb{R}^{n}$ defined as (see [2])
@@
Note that zonotopes are a special case of CZs, where constraints are empty [16], i.e., $A=[\leavevmode\nobreak\ ]$ and $B=[\leavevmode\nobreak\ ]$.
@@
Given an offset $c\!\in\!\mathbb{R}^{n}$, generator $G\!\in\!\mathbb{R}^{n\times h}$, exponent $E\in\mathbb{N}_{0}^{p\times h}$, and constraints $A\in\mathbb{R}^{n_{c}\times q}$, $b\in\mathbb{R}^{n_{c}}$, and $R\in\mathbb{N}_{0}^{p\times q}$, a constrained polynomial zonotope (CPZ) is a compact set in $\mathbb{R}^{n}$ defined as (see [16])
$\mathcal{P}=\bigg{\{}c+\sum_{i=1}^{h}\bigg{(}\prod_{k=1}^{p}\alpha_{(k)}^{E_{(k,i)}}\bigg{)}G_{(\cdot,i)}\leavevmode\nobreak\ \bigg{|}\\ \sum_{i=1}^{q}\bigg{(}\prod_{k=1}^{p}\alpha_{(k)}^{R_{(k,i)}}\bigg{)}A_{(\cdot,i)}=b,\alpha_{(k)}\in[\shortminus 1,1]\bigg{\}}.$ (4)
@@
Given an offset $C\in\mathbb{R}^{m\times n}$, generators $G=\begin{bmatrix}G_{(1)}\!\dots\!G_{(\gamma)}\end{bmatrix}\in\mathbb{R}^{m\times(n\gamma)}$, constraints $A=\begin{bmatrix}A_{(1)}\!\dots\!A_{(\gamma)}\end{bmatrix}\in\mathbb{R}^{n_{c}\times(n_{a}\gamma)}$, and $B\in\mathbb{R}^{n_{c}\times n_{a}}$, a constrained matrix zonotope (CMZ) is a compact set in $\mathbb{R}^{m\times n}$ defined as (see [3])
@@
Note that matrix zonotopes are a special case of CMZs, denoted by $\mathcal{N}=\langle C,G,[\leavevmode\nobreak\ ],[\leavevmode\nobreak\ ],\mathsf{id}\rangle_{\text{CMZ}}$.
@@
Proposition 1 (mergeID [18])
@@
with $\overline{\mathsf{id}}=\begin{bmatrix}\mathsf{id}_{1}&\mathsf{id}_{2}^{(\cdot,\mathcal{H})}\end{bmatrix}$, =$\mathcal{H}=\left\{i\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathsf{id}_{2}^{(i)}\not\in\mathsf{id}_{1}\right\}$, and
@@
$\overline{E}_{2}^{(i,\cdot)}$ =$=\begin{cases}E_{2}^{(j,\cdot)},&\mathrm{if}\leavevmode\nobreak\ \exists j\leavevmode\nobreak\ \overline{\mathsf{id}}_{(i)}=\mathsf{id}_{2}^{(j)}\\ 0_{1\times h_{2}},&\mathrm{otherwise}\end{cases}$
$\overline{R}_{2}^{(i,\cdot)}$ =$=\begin{cases}R_{2}^{(j,\cdot)},&\mathrm{if}\leavevmode\nobreak\ \exists j\leavevmode\nobreak\ \overline{\mathsf{id}}_{(i)}=\mathsf{id}_{2}^{(j)}\\ 0_{1\times q_{2}},&\mathrm{otherwise}\end{cases}$
@@
${\mathcal{P}}_{2}=\Bigg{\{}\begin{bmatrix}3\\ 3\\ 4\end{bmatrix}+\begin{bmatrix}2\\ 3\\ 1\end{bmatrix}\alpha_{(1)}^{3}\alpha_{(2)}^{3}+\begin{bmatrix}2\\ 0\\ 4\end{bmatrix}\alpha_{(1)}^{2}\Biggm{|}\\ \alpha_{(1)}^{2}\alpha_{(2)}^{2}\begin{bmatrix}1\\ 2\end{bmatrix}+\alpha_{(2)}^{3}\begin{bmatrix}3\\ 4\end{bmatrix}=\begin{bmatrix}2\\ 5\end{bmatrix},\alpha_{(1)},\alpha_{(2)}\in[0,1]\Bigg{\}}$ (9)
@@
To preserve dependencies between generators during the addition of two CPZs and prevent over-approximation, we adopt the exact addition operation proposed by [18]. By retaining dependency information, the resulting sets are less conservative, improving the computational efficiency and reliability of the resulting CPZ.
@@ II-D Assumptions
$\begin{array}[]{c}X_{0}^{+}=\begin{bmatrix}x_{(1)}&x_{(2)}&\dots&x_{(T_{0})}\end{bmatrix},\\[2.0pt] X_{0}^{-}=\begin{bmatrix}x_{(0)}&x_{(1)}&\dots&x_{(T_{0}-1)}\end{bmatrix},\\[2.0pt] U_{0}^{-}=\begin{bmatrix}u_{(0)}&u_{(1)}&\dots&u_{(T_{0}-1)}\end{bmatrix}.\end{array}$ (11)
@@
where $C_{\mathcal{M}_{w}}\in\mathbb{R}^{n_{x}\times n_{T}}$ and $G_{\mathcal{M}_{w}}\in\mathbb{R}^{n_{x}\times\gamma_{\mathcal{Z}_{w}}n_{T}}$ with $\mathcal{M}_{w}$ denoting the MZ resulting from the concatenation of multiple noise zonotopes $\mathcal{Z}_{w}$ [3]. Finally, let all system matrices $\begin{bmatrix}\Phi&\Gamma\end{bmatrix}$ of (1) consistent with the data $D_{-}$ be
@@ III Data-driven Reachability Analysis with Non-convex Set Representations
Given system (1) with unknown matrices $\Phi_{\text{tr}}$ and $\Gamma_{\text{tr}}$, the presence of noise implies multiple feasible matrices $\begin{bmatrix}\Phi&\Gamma\end{bmatrix}$ are consistent with the observed data. For the reachability analysis to be reliable, one must incorporate all such data-consistent models. To this end, we extend the offline approach proposed by [3] to continue refining the estimated set of models by leveraging online data.
@@ III-B Intersection Between CMZs
$\hat{A}=\left[\begin{array}[]{cc}\hat{A}_{1}&0_{n_{c_{1}}n_{a_{1}}\times\gamma_{2}}\\ 0_{n_{c_{2}}n_{a_{2}}\times\gamma_{1}}&\hat{A}_{2}\\ \hat{G}_{1}&-\hat{G}_{2}\end{array}\right],\leavevmode\nobreak\ \hat{B}=\left[\begin{array}[]{c}\text{vec}\big{(}{B}_{1}\big{)}\\ \text{vec}\big{(}{B}_{2}\big{)}\\ \text{vec}\big{(}{C}_{2}-{C}_{1}\big{)}\end{array}\right]$
@@
Although the proof is non-trivial, it follows lines of arguments similar to those of [2, Proposition 1]. It is therefore omitted due to space restrictions.
Complexity: The computations of $\text{vec}(B_{1})$, $\text{vec}(B_{2})$, and $\text{vec}(C_{2}-C_{1})$ require $\mathcal{O}(n_{c_{1}}n_{a_{1}})$, $\mathcal{O}(n_{c_{2}}n_{a_{2}})$, and $\mathcal{O}(2n_{x}n)$ operations, respectively. The constructions of $\hat{A}_{1}$, $\hat{A}_{2}$, $\hat{G}_{1}$, and $\hat{G}_{2}$ involve complexities $\mathcal{O}(n_{c_{1}}n_{a_{1}}\gamma_{1})$, $\mathcal{O}(n_{c_{2}}n_{a_{2}}\gamma_{2})$, $\mathcal{O}(n_{x}n\gamma_{1})$, and $\mathcal{O}(n_{x}n\gamma_{2})$, respectively. Thus, the overall computational complexity of intersection (2) is $\mathcal{O}\left(\gamma_{1}(n_{c_{1}}n_{a_{1}}+n_{x}n)+\gamma_{2}(n_{c_{2}}n_{a_{2}}+n_{x}n)\right).$diamond
@@
Accordingly, the reshaped constraint matrix $\hat{A}$ lies in the space $\hat{A}\in\mathbb{R}^{\hat{n}_{{c}}\times\left((\hat{m}/\hat{n}_{c})\cdot(\gamma_{1}+\gamma_{2})\right)},$ where $\hat{A}=\big{[}{\texttt{Convert}}(\hat{A}^{(1)}),\dots,{\texttt{Convert}}(\hat{A}^{(\gamma_{1}+\gamma_{2})})\big{]}.$lrcorner

### v2

@@ I Introduction
Data-driven approaches for reachability can be categorized into deterministic frameworks with bounded uncertainties and stochastic frameworks with probabilistic uncertainty models. Within the deterministic framework, several set representations have been employed to obtain overapproximated reachable sets. For instance, [alanwar2021data] employs matrix zonotopes for representing the set of possible models consistent with noisy data, which is then used to approximate the reachable set as a constrained zonotope, a set representation introduced by [scott2016constrained]. To incorporate prior knowledge and reduce conservatism, [Alanwar2023Datadriven] employs constrained matrix zonotopes and extends the method to polynomial systems and Lipschitz nonlinear systems. Similarly, [park2024data] uses linear time-varying approximations with bounded error terms for nonlinear systems and proposes an ellipsoidal over-approximation of the reachable set by solving a convex optimization problem. For hybrid or logical systems, [siefert2022robust] and [alanwar2025polynomial] introduce hybrid zonotopes and logical polynomial zonotopes, respectively, for representing reachable sets.
@@
Within the stochastic framework, [thorpe2019model, thorpe2021sreachtools] use kernel distribution embeddings in reproducing kernel Hilbert spaces for data-driven reachability, while [devonport2023data] apply empirical inverse Christoffel functions with probably approximately correct (PAC) guarantees, and [griffioen2023data] employ Gaussian process state space models for exact probability measures of trajectories. For human-in-the-loop systems, [choi2023data] and [govindarajan2017data] use Gaussian mixture models and optimized disturbance bounds, respectively. Recent advances also leverage deep learning and statistical methods, e.g., [sivaramakrishnan2024stochastic] approximate stochastic reachability via neural networks, and [hashemi2024statistical] address distribution shift in stochastic cyber-physical systems.
@@
Deterministic approaches to data-driven reachability are preferred when constraint violations are unacceptable, e.g., collision avoidance. This is why regulatory frameworks for safety-critical systems (e.g., DO-178C [rierson2017developing]: Software Considerations in Airborne Systems and Equipment Certification) often require deterministic guarantees rather than probabilistic ones. Furthermore, when data is limited or uncertainty distributions are unknown, deterministic approaches avoid potentially incorrect distributional assumptions, leading to robust guarantees even against worst-case scenarios.
@@
Handling non-convex sets in deterministic reachability is challenging, e.g., obstacle-induced initial sets. While convex sets have a computational advantage, they sacrifice accuracy by overapproximating the reachable set [Alanwar2023Datadriven]. Recent works [dietrich2024nonconvex, wang2023data, djeumou2022fly] have addressed non-convex reachable sets to some degree. However, accurately representing and propagating non-convex sets without introducing excessive conservatism or computational complexity remains an open problem.
@@
The challenge of non-convexity in data-driven reachability is particularly evident in linear systems where one obtains a reachable set by multiplying a set of system matrices consistent with data with a set of possible states. For instance, in [Alanwar2023Datadriven], reachable sets and a set of possible system matrices consistent with data are described using constrained zonotopes (CZs) and constrained matrix zonotopes (CMZs), respectively. However, CZs are not closed under multiplication with CMZs. This is the reason why convex set-based approaches to reachability typically resort to approximations that overbound the resulting set, leading to conservatism.
@@
In this paper, we address this challenge in deterministic data-driven reachability by employing constrained polynomial zonotopes (CPZs) recently introduced by [kochdumper2023constrained]. While [luo2023reachability] presents an exact multiplication for matrix zonotopes (MZs) and polynomial zonotopes (PZs), it does not incorporate constraints into the multiplication itself. Therefore, our key contribution is to add the constraints of the sets in the exact multiplication operation by leveraging the closure property of CPZs when multiplied with CMZs. This extension preserves the non-convex geometry without introducing conservatism, thereby enabling more accurate safety verification for systems that can be more precisely characterized as CMZs—as opposed to just MZs—when dealing with non-convex reachable sets.
@@ II-B Problem Statement
Our goal is to estimate the exact reachable set of the system (1) using input-state data, without explicit knowledge of the system matrices $\Phi_{\text{tr}}$ and $\Gamma_{\text{tr}}$. Moreover, we are interested in the case where the initial set $\mathcal{X}_{0}$ is non-convex, and we must compute a reachable set that preserves this non-convex geometry without resorting to convex relaxations.
@@ II-C Set Representations
Given an offset $c\!\in\!\mathbb{R}^{n}$, generator $G\!\in\!\mathbb{R}^{n\times h}$, and constraints $A\in\mathbb{R}^{n_{c}\times h}$ and $b\in\mathbb{R}^{n_{c}}$, a constrained zonotope (CZ) is a compact set in $\mathbb{R}^{n}$ defined as (see [scott2016constrained])
@@
Note that zonotopes are a special case of CZs, where constraints are empty [kochdumper2023constrained], i.e., $A=[\penalty 10000\ ]$ and $B=[\penalty 10000\ ]$.
@@
Given an offset $c\!\in\!\mathbb{R}^{n}$, generator $G\!\in\!\mathbb{R}^{n\times h}$, exponent $E\in\mathbb{N}_{0}^{p\times h}$, and constraints $A\in\mathbb{R}^{n_{c}\times q}$, $b\in\mathbb{R}^{n_{c}}$, and $R\in\mathbb{N}_{0}^{p\times q}$, a constrained polynomial zonotope (CPZ) is a compact set in $\mathbb{R}^{n}$ defined as (see [kochdumper2023constrained])
$\mathcal{P}=\bigg\{c+\sum_{i=1}^{h}\bigg(\prod_{k=1}^{p}\alpha_{(k)}^{E_{(k,i)}}\bigg)G_{(\cdot,i)}\penalty 10000\ \bigg|\\ \sum_{i=1}^{q}\bigg(\prod_{k=1}^{p}\alpha_{(k)}^{R_{(k,i)}}\bigg)A_{(\cdot,i)}=b,\alpha_{(k)}\in[\shortminus 1,1]\bigg\}.$ (4)
@@
Given an offset $C\in\mathbb{R}^{m\times n}$, generators $G=\begin{bmatrix}G_{(1)}\!\dots\!G_{(\gamma)}\end{bmatrix}\in\mathbb{R}^{m\times(n\gamma)}$, constraints $A=\begin{bmatrix}A_{(1)}\!\dots\!A_{(\gamma)}\end{bmatrix}\in\mathbb{R}^{n_{c}\times(n_{a}\gamma)}$, and $B\in\mathbb{R}^{n_{c}\times n_{a}}$, a constrained matrix zonotope (CMZ) is a compact set in $\mathbb{R}^{m\times n}$ defined as (see [Alanwar2023Datadriven])
@@
Note that matrix zonotopes are a special case of CMZs, denoted by $\mathcal{N}=\langle C,G,[\penalty 10000\ ],[\penalty 10000\ ],\mathsf{id}\rangle_{\text{CMZ}}$.
@@
Proposition 1 (mergeID [kochdumper2020sparse])
@@
with $\overline{\mathsf{id}}=\begin{bmatrix}\mathsf{id}_{1}&\mathsf{id}_{2}^{(\cdot,\mathcal{H})}\end{bmatrix}$, $\mathcal{H}=\left\{i\penalty 10000\ |\penalty 10000\ \mathsf{id}_{2}^{(i)}\not\in\mathsf{id}_{1}\right\}$, and
@@
$\overline{E}_{2}^{(i,\cdot)}$ $=\begin{cases}E_{2}^{(j,\cdot)},&\mathrm{if}\penalty 10000\ \exists j\penalty 10000\ \overline{\mathsf{id}}_{(i)}=\mathsf{id}_{2}^{(j)}\\ 0_{1\times h_{2}},&\mathrm{otherwise}\end{cases}$
$\overline{R}_{2}^{(i,\cdot)}$ $=\begin{cases}R_{2}^{(j,\cdot)},&\mathrm{if}\penalty 10000\ \exists j\penalty 10000\ \overline{\mathsf{id}}_{(i)}=\mathsf{id}_{2}^{(j)}\\ 0_{1\times q_{2}},&\mathrm{otherwise}\end{cases}$
@@
${\mathcal{P}}_{2}=\Bigg\{\begin{bmatrix}3\\ 3\\ 4\end{bmatrix}+\begin{bmatrix}2\\ 3\\ 1\end{bmatrix}\alpha_{(1)}^{3}\alpha_{(2)}^{3}+\begin{bmatrix}2\\ 0\\ 4\end{bmatrix}\alpha_{(1)}^{2}\Biggm|\\ \alpha_{(1)}^{2}\alpha_{(2)}^{2}\begin{bmatrix}1\\ 2\end{bmatrix}+\alpha_{(2)}^{3}\begin{bmatrix}3\\ 4\end{bmatrix}=\begin{bmatrix}2\\ 5\end{bmatrix},\alpha_{(1)},\alpha_{(2)}\in[-1,1]\Bigg\}$ (9)
@@
To preserve dependencies between generators during the addition of two CPZs and prevent over-approximation, we adopt the exact addition operation proposed by [kochdumper2020sparse]. By retaining dependency information, the resulting sets are less conservative, improving the computational efficiency and reliability of the resulting CPZ.
@@ II-D Assumptions
[…]
@@
where $C_{\mathcal{M}_{w}}\in\mathbb{R}^{n_{x}\times n_{T}}$ and $G_{\mathcal{M}_{w}}\in\mathbb{R}^{n_{x}\times\gamma_{\mathcal{Z}_{w}}n_{T}}$ with $\mathcal{M}_{w}$ denoting the MZ resulting from the concatenation of multiple noise zonotopes $\mathcal{Z}_{w}$ [Alanwar2023Datadriven]. Finally, let all system matrices $\begin{bmatrix}\Phi&\Gamma\end{bmatrix}$ of (1) consistent with the data $D_{-}$ be
@@ III Data-driven Reachability Analysis with Non-convex Set Representations
Given system (1) with unknown matrices $\Phi_{\text{tr}}$ and $\Gamma_{\text{tr}}$, the presence of noise implies multiple feasible matrices $\begin{bmatrix}\Phi&\Gamma\end{bmatrix}$ are consistent with the observed data. For the reachability analysis to be reliable, one must incorporate all such data-consistent models. To this end, we extend the offline approach proposed by [Alanwar2023Datadriven] to continue refining the estimated set of models by leveraging online data.
@@ III-B Intersection Between CMZs
$\hat{A}=\left[\begin{array}[]{cc}\hat{A}_{1}&0_{n_{c_{1}}n_{a_{1}}\times\gamma_{2}}\\ 0_{n_{c_{2}}n_{a_{2}}\times\gamma_{1}}&\hat{A}_{2}\\ \hat{G}_{1}&-\hat{G}_{2}\end{array}\right],\penalty 10000\ \hat{B}=\left[\begin{array}[]{c}\text{vec}\big({B}_{1}\big)\\ \text{vec}\big({B}_{2}\big)\\ \text{vec}\big({C}_{2}-{C}_{1}\big)\end{array}\right]$
@@
Although the proof is non-trivial, it follows lines of arguments similar to those of [scott2016constrained, Proposition 1]. It is therefore omitted due to space restrictions.
Complexity: The computations of $\text{vec}(B_{1})$, $\text{vec}(B_{2})$, and $\text{vec}(C_{2}-C_{1})$ require $\mathcal{O}(n_{c_{1}}n_{a_{1}})$, $\mathcal{O}(n_{c_{2}}n_{a_{2}})$, and $\mathcal{O}(2n_{x}n)$ operations, respectively. The constructions of $\hat{A}_{1}$, $\hat{A}_{2}$, $\hat{G}_{1}$, and $\hat{G}_{2}$ involve complexities $\mathcal{O}(n_{c_{1}}n_{a_{1}}\gamma_{1})$, $\mathcal{O}(n_{c_{2}}n_{a_{2}}\gamma_{2})$, $\mathcal{O}(n_{x}n\gamma_{1})$, and $\mathcal{O}(n_{x}n\gamma_{2})$, respectively. Thus, the overall computational complexity of intersection (2) is $\mathcal{O}\left(\gamma_{1}(n_{c_{1}}n_{a_{1}}+n_{x}n)+\gamma_{2}(n_{c_{2}}n_{a_{2}}+n_{x}n)\right).$ $\diamond$
@@
[…]